# Investigating Length Issues in Document-level Machine Translation

### v1

@@ Abstract
Transformer architectures are increasingly effective at processing and generating very long chunks of texts, opening new perspectives for document-level machine translation (MT). In this work, we challenge the ability of MT systems to handle texts comprising up to several thousands of tokens. We design and implement a new approach designed to precisely measure the effect of length increments on MT outputs. Our experiments with two representative architectures unambiguously show that (a) translation performance decreases with the length of the input text; (b) the position of sentences within the document matters and translation quality is higher for sentences occurring earlier in a document. We further show that manipulating the distribution of document lengths and of positional embeddings only marginally mitigates such problems. Our results suggest that even though document-level MT is computationally feasible, it does not yet match the performance of sentence-based MT.
@@ 1 Introduction
The ability of today’s neural MT models—relying on encoder-decoder or decoder-only architectures—to handle large context lengths, up to thousands of tokens (Peng et al., 2024), opens new perspectives to go beyond context-augmented MT and develop fully-fledged document-level MT, where the entire document context is available at once, and where the target text is generated in a single pass. [^1] These perspectives are, for instance, explored in the latest edition of the WMT shared task on General Machine Translation (Kocmi et al., 2024). Two main technical novelties have made this possible: (a) more efficient computation in the attention layers (Tay et al., 2023a) and (b) changes in the design of positional encodings (PEs). In particular, replacing sinusoidal absolute PEs (APEs) of (Vaswani et al., 2017) with methods like ALIBI Press et al. (2022) and RoPE (Su et al., 2024), which lend themselves well to length extrapolation (Sun et al., 2023; Zhao et al., 2024) seem to make today’s transformers amenable to the processing of arbitrarily long contexts (Mohtashami and Jaggi, 2023; Han et al., 2024).
In this work, we challenge the ability of contemporary MT models to effectively handle long spans of texts. For this, we develop a new methodology for assessing the impact of length variations on MT performance. We perform a series of controlled experiments with two representative neural MT systems, where the same documents are processed by chunks of increasing lengths in a document-level manner and show that (a) MT performance tends to degrade with the length of the source document; (b) length issues happen even for in-distribution lengths and get worse when extrapolating to unseen document lengths, and (c) most of the degradation happens in the final parts of the translation. Hypothesising that this may be due to a mismatch between the distribution of train and test PEs, we explore a possible mitigation, which flattens the distribution of PEs during training. We observe a consistent improvement of automatic metric scores for the APE-based vanilla encoder-decoder model, while the RoPE based decoder-only model remains mostly unaffected. In summary, our main contributions are: (a) a new approach to the detection and diagnosis of length issues in document-level MT; (b) a new variant of the SHAPE Kiyono et al. (2021) method, which improves the distribution of PEs during training; (c) a confirmation that (perhaps for lack of appropriate document-level evaluation tool) sentence-level MT remains a strong baseline in most settings.
@@ 2.1 Document-level MT
Translating sentence by sentence, even with augmented contexts, still fails to capture coherence and consistency issues (Fernandes et al., 2023), motivating Doc2Doc scenarios to process documents as a whole. This can be done with concatenation-based methods (Tiedemann and Scherrer, 2017; Sun et al., 2022; Karpinska and Iyyer, 2023), along with sliding window attention (Zhuocheng et al., 2023; Liu et al., 2023) or group attention (Bao et al., 2021) to address the quadratic complexity. Other strategies focus on improving training through data augmentation with a balanced length distribution (Sun et al., 2022) or richer context-dependent phenomena (Lupo et al., 2022a; Wu et al., 2024); or better training strategies with multilingual denoising pretraining (Lee et al., 2022), adapted loss functions (Lupo et al., 2022b), or enriched positional encodings (Li et al., 2023; Lupo et al., 2023). Multiple methods have recently emerged for large language models (LLMs) (Wang et al., 2023), such as the two-stage training recipe with the use of a monolingual corpus and high-quality parallel documents (Xu et al., 2024; Alves et al., 2024), or applying LLMs as post-editors (Koneru et al., 2024).
@@ 2.2 Extrapolating PEs
Since the attention mechanism is position-agnostic, PEs are used to provide position information in Transformer models. PEs embed the absolute token position (APEs) (Vaswani et al., 2017), or the relative distance between tokens (RPEs) (Shaw et al., 2018; Raffel et al., 2020; Press et al., 2022), with RoPE (Su et al., 2024) being the go-to approach in recent LLMs such as Llama2 (Touvron et al., 2023). Despite RPEs yield better length extrapolation ability than APEs, both of them struggle to efficiently extrapolate input lengths beyond the predefined maximum training length (Dai et al., 2019; Chen et al., 2023; Peng et al., 2024; Zhao et al., 2024), motivating the development of input extension methods for PEs.
For APEs, SHAPE (Kiyono et al., 2021) offsets all indices in a sequence by some random values. Its authors show that this simple technique mimics the computation of RPEs at a much smaller cost and helps to improve the interpolation abilities of a vanilla encoder-decoder model, as measured by BLEU (Papineni et al., 2002) scores with long pseudo-documents. Our experiments confirm that this technique is effective using actual document contexts and a sounder experimental methodology, based on paired tests, and using COMET (Rei et al., 2020).disabledisabletodo: disableRead this one ! Sinha et al.’s experiments (2022) adopt a setting similar to ours, offsetting the absolute value of APEs’ input to evaluate their ability to capture relative distances between tokens. Their results, like ours, illustrate the lack of robustness of APEs and suggest that they overfit their training data.
@@ 3.1 Holistic Document-Level MT
    in Doc2Doc, input texts are longer, causing a computational overhead due to the quadratic complexity of attention (Tay et al., 2023b).
@@
    for longer inputs, attention weights are spread over a larger number of tokens (Herold and Ney, 2023); yet, at each decoding step, most attention needs to remain concentrated on the corresponding local source context (Bao et al., 2021). This is in contrast with Doc2Sent, where sentence alignment is readily available.
@@
    decoding longer sequences increases the impact of search errors and of exposure bias (Ranzato et al., 2016). The beam search also becomes more difficult due to the input length.
@@ 3.2 Shades of BLEU
Answering such questions requires metrics for comparing holistic translations with sentence-based translations: as the number of segments produced by the former may differ from the number of source segments, a basic requirement is that they allow the evaluation of translation hypotheses with more (or fewer) sentences than the source (for quality estimation scores) and/or the reference (for reference-based metrics). Yet, most existing document-level MT approaches still rely on BLEU scores Papineni et al. (2002), despite its well-documented shortcomings (Callison-Burch et al., 2006; Reiter, 2018; Mathur et al., 2020); or rather a variant dubbed d-BLEU by Liu et al. (2020). [^2] Hendy et al. (2023) also consider a variant of COMET Rei et al. (2022) while Zhuocheng et al. (2023) introduce d-ChrF, a document-level version of ChrF (Popović, 2015). We accordingly focus on BLEU in this section, noting that the same questions would need to be addressed with any metric relying on sentence-based surface comparison (e.g., METEOR Banerjee and Lavie (2005), TER Snover et al. (2006), BertSCORE (Zhang et al., 2020), PRISM (Thompson and Post, 2020), COMET (Rei et al., 2020), and many others).
BLEU is computed by counting, sentence by sentence, the number of $n$-grams (for $n\in[1\!:\!4]$) shared by each translation hypothesis and its human reference. These counts are aggregated and turned into frequencies, then averaged (geometrically) at the corpus level. Finally, a length penalty is applied to degrade the score when the cumulated length of the hypotheses is shorter than that of the references. BLEU is a corpus-level score that depends on sentence alignments. d-BLEU is also a global score but counts common $n$-grams at the document level. This implies that d-BLEU, which records matches for larger spans than BLEU, delivers higher scores, as the opportunities to match $n$-grams are greater for a wider window. [^3] This effect is well known, e.g. in (Koehn and Knowles, 2017, Figure 7), where BLEU increases when considering sentence groups of increasing lengths (at least for a certain length range), where we would expect a decrease, as the length is often linked to syntactic complexity and therefore to translation difficulty. We reproduce this observation in Figure 2. These two scores cannot be compared, and we contend that their shortcomings make them inappropriate for analysing length-related issues in MT.
@@
Figure 1: Top: probability of observing training position $i$ ($P(i)$) for a sentence of length $l=200$, with the standard training ($k_{i}=0$, left) and with our uniform sampling scheme (right) for $M=512$. Middle : original and unifPE $P(i)$ for training set TED-G and $M=2048$. Bottom: original and unifPE $P(i)$ for TED-U and $M=2048$.
@@ 3.3 Evaluating Length Issues in MT
Another recurring methodological caveat with the length-related evaluation is related to the way scores are compared. For instance, in (Sun et al., 2022, Fig. 1) BLEU scores are reported for buckets of sentences of varying lengths in a plot which suggests that performance increases with length (up to a certain extent). Such vizualisations are misleading, as global BLEU scores should only be compared when measured with the same corpus.
@@
In our experiments, we consider two ways of designing $\mathcal{S}$: (a) at the document level, where each $s_{i}$ is a document and the evaluation is the ds-BLEU score introduced in Section 3.2; [^4] We use SacreBLEU (Post, 2018) with signature:
@@ 4.2 Uniform SHAPEs (unifPE)
disabledisabletodo: disable1. Goal, 2. Intuition, 3. A sound solution
We assume a training set of texts $s_{1}\dots s_{N}$ of respective lengths $l_{1}\dots l_{N}$, and a maximum model length of $M$, with $\forall i,M>l_{i}$. Training with text $s_{i}$ creates training samples for positions $i$ in $[1\!:\!l_{i}]$. For the whole corpus, positions from $1$ (observed $N$ times) to $l_{max}=\mathrm{max}_{i=1\dots N}(l_{i})$ are observed, with larger indexes being less trained than smaller ones. Positions indices in $[l_{max}\!:\!M]$ are never observed. We wish to make the training PE distribution more even, so that all positions in $[1\!:\!M]$ are equally well-trained, which should also help to extrapolate PEs for indexes larger than $l_{max}$.
@@
disabledisabletodo: disableperhaps mention SHAPE
This approach is reminiscent of SHAPE (Kiyono et al., 2021); while SHAPE chooses the offset $k_{i}$ uniformly at random in a fixed interval to simulate relative PEs, we sample $k_{i}$ non-uniformly to ensure that all indexes are equally represented in training.
@@ 5.1 Datasets
NLLB FT-U Unif-U FT-G Unif-G
sent-256 9.2 (0.00) 0.8 (0.00) -2.1 (0.00) 0.4 (0.03) -0.7 (0.00)
256-512 19.1 (0.00) - -0.4 (0.03) - -
512-768 6.9 (0.00) - -0.6 (0.00) 5.9 (0.00) 2.5 (0.00)
768-1024 - 0.5 (0.01) - 7.2 (0.00) 3.0 (0.00)
1024-1200 2.2 (0.00) 3.5 (0.00) 1.9 (0.00) 4.0 (0.00) 3.3 (0.00)
1200-1600 - 6.8 (0.00) 6.5 (0.00) 5.8 (0.00) 5.5 (0.00)
1600-2048 1.9 (0.00) 5.2 (0.00) 4.5 (0.00) 4.2 (0.00) 5.7 (0.00)
sent-256 16.7 (0.00) 3.5 (0.00) 1.7 (0.00) 2.7 (0.00) 2.1 (0.00)
256-512 20.7 (0.00) - -0.4 (0.03) 0.6 (0.00) -
512-768 5.6 (0.00) - - 11.6 (0.00) 7.2 (0.00)
768-1024 5.2 (0.00) 2.3 (0.00) 0.8 (0.04) 10.4 (0.00) 7.6 (0.00)
1024-1200 - 7.4 (0.00) 4.3 (0.00) 3.8 (0.00) 4.7 (0.00)
1200-1600 6.1 (0.0) 9.6 (0.00) 13.4 (0.00) 5.4 (0.00) 5.5 (0.00)
1600-2048 - 5.1 (0.00) 5.0 (0.00) 3.9 (0.00) 5.6 (0.00)
Tower FT-U Unif-U FT-G Unif-G
sent-256 - - - - -
256-512 - 0.9 (0.02) 0.8 (0.05) 0.6 (0.00) 0.6 (0.00)
512-768 - - - 0.6 (0.04) -
768-1024 3.4 (0.00) - 1.0 (0.05) 1.7 (0.01) 1.2 (0.01)
1024-1200 - - - - -
1200-1600 4.7 (0.00) 1.7 (0.01) 2.1 (0.01) 1.6 (0.02) 2.0 (0.01)
1600-2048 5.9 (0.00) 7.5 (0.00) 6.5 (0.00) 8.1 (0.00) 7.8 (0.00)
sent-256 3.9 (0.00) 2.3 (0.00) 2.4 (0.00) 2.3 (0.00) 2.3 (0.00)
256-512 - 0.4 (0.02) - 0.2 (0.03) 0.3 (0.02)
512-768 - - - 0.3 (0.00) -
768-1024 2.9 (0.00) 1.0 (0.04) 0.5 (0.00) 1.1 (0.03) 0.8 (0.00)
1024-1200 - - 1.0 (0.03) - -
1200-1600 6.2 (0.00) 1.7 (0.00) 1.8 (0.02) - 1.9 (0.01)
1600-2048 8.7 (0.00) 10.0 (0.00) 8.9 (0.00) 11.0 (0.00) 10.2 (0.00)
Table 3: Average differences (and p-values) evaluated on ds-BLEU (top) of full TED talks and on 100×100$\times$100 ×COMET (bottom) of realigned parallel sentences, for NLLB (left) and TowerBase (right) models. - for p-values $>0.05$. U and G respectively denote TED-U and TED-G.
@@
Our training set consists of pseudo-documents from both the training and validation splits of IWSLT-2016.999https://wit3.fbk.eu/2016-01 Our goal is to simulate real corpora of parallel documents with source documents shorter than a certain length $l_{max}$ – using $l_{max}=1024$. We split all document pairs whose source side is longer than 1024 tokensdisabledisabletodo: disabletokens? into fragments. [^10] All the statistics in tokens is counted using the tokeniser of NLLB (Costa-jussà et al., 2024). For each document pair, we iterate the following procedure: (1) sample a maximum pseudo-document length $l^{\prime}_{i}$ following the same Gaussian-like length distribution as the full TED talks with $l^{\prime}_{i}<l_{max}$; (2) concatenate consecutive sentence pairs up to $l^{\prime}_{i}$ to form a training pseudo-documents $s_{i}$. The resulting distribution of document lengths is displayed in Figure 3. The development set is built similarly, using document pairs from IWSLT tst2010 and tst2011. We denote these training datasets as TED-G (G for Gaussian). As discussed in Section 4, we consider another dataset generation strategy, which produces a more balanced length distribution, for which we do as above but we sample uniformly: $l^{\prime}_{i}\sim U(128,l_{max})$. [^11] Short pseudo-documents continue to be slightly over-represented, because the last pseudo-document in any given talk is often strictly shorter than the desired length $l^{\prime}_{i}$. Fine-tuning with the resulting TED-U corpus allows us to contrast two distributions with differences in document length.
@@
To evaluate MT systems for their ability to handle documents of varying sizes and extrapolate beyond the training samples, we build a series of test sets of increasing document lengths. For each document in IWSLT tst2014, tst2015, tst2016 and tst2017, we accumulate consecutive sentence pairs into parallel pseudo-documents such that all resulting source texts have a length close to $l_{max}$, with $l_{max}\in\{256,512,1024,1200,1600,2048\}$. [^12] At the end of each talk, we concatenate the last parallel sentences into the last pseudo-document if they are shorter than $50$ to avoid exceedingly short parallel sequences. Contrarily to training sets, test sets are homogeneous in length. Statistics are in Table 1 with more details in Table 8. Evaluation is always performed with complete original talks, after concatenating and aligning all the corresponding parts.

### v2

@@ Abstract
Transformer architectures are increasingly effective at processing and generating very long chunks of texts, opening new perspectives for document-level machine translation (MT). In this work, we challenge the ability of MT systems to handle texts comprising up to several thousands of tokens. We design and implement a new approach designed to precisely measure the effect of length increments on MT outputs. Our experiments with two representative architectures unambiguously show that (a) translation performance decreases with the length of the input text; (b) the position of sentences within the document matters, and translation quality is higher for sentences occurring earlier in a document. We further show that manipulating the distribution of document lengths and of positional embeddings only marginally mitigates such problems. Our results suggest that even though document-level MT is computationally feasible, it does not yet match the performance of sentence-based MT.
@@
footnotetext: ©  2025 The authors. This article is licensed under a Creative Commons 4.0 licence, no derivative works, attribution, CC-BY-ND.
@@ 1 Introduction
The ability of today’s neural MT models—relying on encoder-decoder or decoder-only architectures—to handle large context lengths, up to thousands of tokens (Peng et al., 2024), opens new perspectives to go beyond context-augmented MT and develop fully-fledged document-level MT, where the entire document context is available at once, and where the target text is generated in a single pass. [^1] These perspectives are, for instance, explored in the latest edition of the WMT shared task on General Machine Translation (Kocmi et al., 2024). Two main technical novelties have made this possible: (a) more efficient computation in the attention layers (Tay et al., 2022) and (b) changes in the design of positional encodings (PEs). In particular, replacing the sinusoidal absolute PEs (APEs) of (Vaswani et al., 2017) with methods like ALIBI Press et al. (2022) and RoPE (Su et al., 2024), which lend themselves well to length extrapolation (Sun et al., 2023; Zhao et al., 2024), seems to make today’s transformers amenable to the processing of arbitrarily long contexts (Mohtashami and Jaggi, 2023; Han et al., 2024).
In this work, we challenge the ability of contemporary MT models to effectively handle long spans of texts. For this, we develop a new methodology for assessing the impact of length variations on MT performance. We perform a series of controlled experiments with two representative neural MT systems, where the same documents are processed by chunks of increasing lengths in a document-level manner and show that (a) MT performance tends to degrade with the length of the source document, (b) length issues happen even for in-distribution lengths and get worse when extrapolating to unseen document lengths, and (c) most of the degradation happens in the final parts of the translation. Hypothesising that this may be due to a mismatch between the distribution of train and test PEs, we explore a possible mitigation, which flattens the distribution of PEs during training. We observe a consistent improvement of automatic metric scores for the APE-based vanilla encoder-decoder model, while the RoPE based decoder-only model remains mostly unaffected. In summary, our main contributions are: (a) a new approach to the detection and diagnosis of length issues in document-level MT, (b) a new variant of the SHAPE Kiyono et al. (2021) method, which improves the distribution of PEs during training, and (c) a confirmation that (perhaps for lack of an appropriate document-level evaluation tool) sentence-level MT remains a strong baseline in most settings.
@@ 2.1 Document-level MT
Translating sentence by sentence, even with augmented contexts, still fails to capture phenomena related to coherence and consistency (Fernandes et al., 2023), motivating Doc2Doc approaches to process documents as a whole. This can be done with concatenation-based methods (Tiedemann and Scherrer, 2017; Sun et al., 2022; Karpinska and Iyyer, 2023), along with sliding window attention (Zhuocheng et al., 2023; Liu et al., 2023) and group attention (Bao et al., 2021) to address the issue of quadratic complexity. Other strategies include focusing on improving training through data augmentation with a balanced length distribution (Sun et al., 2022) and richer context-dependent phenomena (Lupo et al., 2022a; Wu et al., 2024), or on better training strategies with multilingual denoising pretraining (Lee et al., 2022), adapted loss functions (Lupo et al., 2022b), and enriched positional encodings (Li et al., 2023; Lupo et al., 2023). Multiple methods have recently emerged for large language models (LLMs) (Wang et al., 2023), which also show a decline in translation quality as input length increases (Wang et al., 2024). [^2] They also confirmed the effectiveness of training LLMs on documents of varied sizes (similar to Sun et al. (2022)). These include a two-stage training recipe with the use of a monolingual corpus and high-quality parallel documents (Xu et al., 2024; Alves et al., 2024), and applying LLMs as post-editors (Koneru et al., 2024).
@@ 2.2 Extrapolating PEs
Since self-attention is position-agnostic, PEs are used to provide position information in Transformer models. PEs embed the absolute token position (APEs) (Vaswani et al., 2017), or the relative distance between tokens (RPEs) (Shaw et al., 2018; Raffel et al., 2020; Press et al., 2022), with RoPE (Su et al., 2024) being the go-to approach in recent LLMs such as Llama2 (Touvron et al., 2023). Despite RPEs yielding better length extrapolation ability than APEs, both of them struggle to efficiently extrapolate input lengths beyond the predefined maximum training length (Dai et al., 2019; Chen et al., 2023; Peng et al., 2024; Zhao et al., 2024), motivating the development of input extension methods for PEs.
For APEs, SHAPE (Kiyono et al., 2021) offsets all indices in a sequence by some random values. Its authors show that this simple technique mimics the computation of RPEs at a much smaller cost and helps to improve the interpolation abilities of a vanilla encoder-decoder model, as measured by BLEU (Papineni et al., 2002) with long pseudo-documents. Our experiments confirm that this technique is effective using actual document contexts and a sounder experimental methodology, based on paired tests, and using COMET (Rei et al., 2020).disabledisabletodo: disableRead this one ! Sinha et al.’s (2022) experiments adopt a setting similar to ours, offsetting the absolute value of APEs’ input to evaluate their ability to capture relative distances between tokens. Their results, like ours, illustrate the lack of robustness of APEs and suggest that they overfit their training data.
@@ 3.1 Holistic Document-Level MT
    in Doc2Doc, input texts are longer, causing a computational overhead due to the quadratic complexity of attention (Tay et al., 2022).
@@
    for longer inputs, attention weights are spread over a larger number of tokens (Herold and Ney, 2023); however, at each decoding step, most attention needs to remain concentrated on the corresponding local source context (Bao et al., 2021). This is in contrast with Doc2Sent, where sentence alignment is readily available.
@@
    decoding longer sequences increases the impact of search errors and of exposure bias (Ranzato et al., 2016). Beam search also becomes more difficult due to the input length.
@@ 3.2 Shades of BLEU
Answering such questions requires metrics for comparing holistic translations with sentence-based translations: as the number of segments produced by the former may differ from the number of source segments, a basic requirement is that they allow the evaluation of translation hypotheses with more (or fewer) sentences than the source (for quality estimation scores) and/or the reference (for reference-based metrics). However, most existing document-level MT approaches still rely on BLEU Papineni et al. (2002), despite its well-documented shortcomings (Callison-Burch et al., 2006; Reiter, 2018; Mathur et al., 2020; Dahan et al., 2024); or rather a variant dubbed d-BLEU by Liu et al. (2020). [^3] Hendy et al. (2023) also consider a variant of COMET Rei et al. (2022) while Zhuocheng et al. (2023) introduce d-ChrF, a document-level version of ChrF (Popović, 2015). We accordingly focus on BLEU in this section, noting that the same questions would need to be addressed with any metric relying on sentence-based surface comparison (e.g., METEOR Banerjee and Lavie (2005), TER Snover et al. (2006), BertSCORE (Zhang et al., 2020), PRISM (Thompson and Post, 2020), COMET (Rei et al., 2020), and many others). [^4] We choose to evaluate using the standard metrics, BLEU and COMET, rather than evaluation approaches specifically designed to test the use of increased context. This choice is motivated by the fact that the score differences we observe reveal a significant degradation in translation quality for longer documents, indicating greater problems than those targeted by finer-grained evaluation techniques.
BLEU is computed by counting, sentence by sentence, the number of $n$-grams (for $n\in[1\!:\!4]$) shared by each translation hypothesis and its human reference. These counts are aggregated and turned into frequencies, then averaged (geometrically) at the corpus level. Finally, a length penalty is applied to degrade the score when the cumulated length of the hypotheses is shorter than that of the references. BLEU is a corpus-level score that depends on sentence alignments. d-BLEU is also a global score but counts common $n$-grams at the document level. As a consequence, d-BLEU, which records matches for larger spans than BLEU, delivers higher scores, as the opportunities to match $n$-grams are greater for a wider window. [^5] This effect is well known, e.g. in (Koehn and Knowles, 2017, Figure 7), where BLEU increases when considering sentence groups of increasing lengths (at least for a certain length range), where we would expect a decrease, as the length is often linked to syntactic complexity and therefore to translation difficulty. We reproduce this observation in Figure 2. These two scores cannot be compared, and we contend that their shortcomings make them inappropriate for analysing length-related issues in MT.
@@
Figure 1: Top: probability of observing training position $i$ ($P(i)$) for a sentence of length $l=200$, with standard training ($k_{i}=0$, left) and with our uniform sampling scheme (right) for $M=512$. Middle: original, unifPE, and SHAPE $P(i)$ for training set TED-G and $M=2048$. Bottom: original, unifPE and SHAPE $P(i)$ for TED-U and $M=2048$.
@@ 3.3 Evaluating Length Issues in MT
Another recurring methodological caveat with length-related evaluation is related to the way scores are compared. For instance, in (Sun et al., 2022, Figure 1) BLEU scores are reported for buckets of sentences of varying lengths in a plot which suggests that performance increases with length (up to a certain extent). Such vizualisations are misleading, as global BLEU scores should only be compared when measured with the same corpus.
@@
In our experiments, we consider two ways of presenting $\mathcal{S}$: (a) at the document level, where each $s_{i}$ is a document and the evaluation is the ds-BLEU score introduced in Section 3.2, [^6] We use SacreBLEU (Post, 2018) with signature:
@@ 4.2 Uniform SHAPEs (unifPE)
We assume a training set of texts $s_{1}\dots s_{N}$ of respective lengths $l_{1}\dots l_{N}$, and a maximum model length of $M$, with $\forall i,M>l_{i}$. Training with text $s_{i}$ creates training samples for positions $i$ in $[1\!:\!l_{i}]$. For the whole corpus, positions from $1$ (observed $N$ times) to $l_{max}=\mathrm{max}_{i=1\dots N}(l_{i})$ are observed, with larger indexes being less trained than smaller ones. Positions indices in $[l_{max}\!:\!M]$ are never observed. We wish to make the training PE distribution more even, so that all positions in $[1\!:\!M]$ are equally well-trained, which should also help to extrapolate PEs for indices larger than $l_{max}$.
@@
This approach is reminiscent of SHAPE (Kiyono et al., 2021); while SHAPE chooses the offset $k_{i}$ uniformly at random in a fixed interval to simulate relative PEs, which reduces the frequency of small position indices, we sample $k_{i}$ non-uniformly to ensure that all indexes are equally represented in training.
@@ 5.1 Datasets
NLLB FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
sent-256 9.2 0.8 -2.1 -2.5 0.4 -0.7 -1.7
256-512 19.1 - -0.4 1.4 - - 2.1
512-768 6.9 - -0.6 - 5.9 2.5 5.3
768-1024 - 0.5 - - 7.2 3.0 3.7
1024-1200 2.2 3.5 1.9 4.1 4.0 3.3 4.1
1200-1600 - 6.8 6.5 5.4 5.8 5.5 4.9
1600-2048 1.9 5.2 4.5 3.1 4.2 5.7 6.0
sent-256 16.7 3.5 1.7 1.3 2.7 2.1 1.9
256-512 20.7 - -0.4 2.4 0.6 - 4.2
512-768 5.6 - - - 11.6 7.2 8.1
768-1024 5.2 2.3 0.8 - 10.4 7.6 7.8
1024-1200 - 7.4 4.3 6.9 3.8 4.7 4.6
1200-1600 6.1 9.6 13.4 8.3 5.4 5.5 5.6
1600-2048 - 5.1 5.0 5.9 3.9 5.6 5.0
Tower FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
sent-256 - - - - - - -
256-512 - 0.9 0.8 0.8 0.6 0.6 0.5
512-768 - - - - 0.6 - -
768-1024 3.4 - 1.0 1.2 1.7 1.2 2.1
1024-1200 - - - - - - -
1200-1600 4.7 1.7 2.1 - 1.6 2.0 1.5
1600-2048 5.9 7.5 6.5 7.3 8.1 7.8 7.3
sent-256 3.9 2.3 2.4 2.3 2.3 2.3 2.2
256-512 - 0.4 - - 0.2 0.3 0.3
512-768 - - - 0.5 0.3 - -
768-1024 2.9 1.0 0.5 - 1.1 0.8 1.2
1024-1200 - - 1.0 0.9 - - -
1200-1600 6.2 1.7 1.8 - - 1.9 1.8
1600-2048 8.7 10.0 8.9 9.1 11.0 10.2 9.2
Table 3: Average differences evaluated on ds-BLEU (top) of full TED talks and on 100×100$\times$100 ×COMET (bottom) of realigned parallel sentences, between translations in increasing context size, for NLLB (left) and TowerBase (right) models. U and G respectively denote TED-U and TED-G. A positive value means that shorter segments result in higher scores than longer ones. Text in olive for p-values $>0.01$. - for p-values $>0.05$.
@@
Our training set consists of pseudo-documents from both the training and validation splits of IWSLT-2016.111111https://wit3.fbk.eu/2016-01 Our goal is to simulate real corpora of parallel documents with source documents shorter than a certain length $l_{max}$ – using $l_{max}=1024$. We split all document pairs whose source side is longer than 1024 tokens into fragments. [^12] All statistics counted in tokens use the tokeniser of NLLB (Costa-jussà et al., 2024). For each document pair, we iterate the following procedure: (1) sample a maximum pseudo-document length $l^{\prime}_{i}$ following the same Gaussian-like length distribution as the full TED talks with $l^{\prime}_{i}<l_{max}$, (2) concatenate consecutive sentence pairs up to $l^{\prime}_{i}$ to form a training pseudo-document $s_{i}$. The resulting distribution of document lengths is displayed in Figure 3 in Appendix A.3. The development set is built similarly, using document pairs from IWSLT tst2010 and tst2011. We denote these training datasets as TED-G (G for Gaussian). As discussed in Section 4, we consider another dataset generation strategy, which produces a more balanced length distribution, for which we do as above but we sample uniformly: $l^{\prime}_{i}\sim U(128,l_{max})$. [^13] Short pseudo-documents continue to be slightly over-represented, because the last pseudo-document in any given talk is often strictly shorter than the desired length $l^{\prime}_{i}$. Fine-tuning with the resulting TED-U corpus allows us to contrast two distributions with differences in document length.
@@
To evaluate MT systems for their ability to handle documents of varying sizes and extrapolate beyond the training samples, we build a series of test sets of increasing document lengths. For each document in IWSLT tst2014, tst2015, tst2016 and tst2017, we accumulate consecutive sentence pairs into parallel pseudo-documents such that all resulting source texts have a length close to $l_{max}$, with $l_{max}\in\{256,512,1024,1200,1600,2048\}$. [^14] At the end of each talk, we concatenate the last parallel sentences into the last pseudo-document if they are shorter than $50$ to avoid exceedingly short parallel sequences. Contrarily to training sets, test sets are homogeneous in length. Statistics are in Table 1 with more details in Appendix A.3. Evaluation is always performed with complete original talks, after concatenating and aligning all the corresponding parts.